# CausalRAG: Integrating Causal Graphs into Retrieval-Augmented Generation

### v1

@@ 1 Introduction
The rapid advancements in large language models (LLMs) have revolutionized the field of natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on pre-trained knowledge limits their ability to integrate and reason over dynamically updated external information, particularly in knowledge intensive domains such as academic research. Retrieval-Augmented Generation (RAG) has emerged as a promising framework to address this limitation Lewis et al. (2021), combining retrieval mechanisms with generative capabilities to enhance contextual understanding and response quality.
@@
Recent research has focused on improving RAG along two primary directions: 1) enhancing retrieval efficiency and integration mechanisms by designing more adaptive and dynamic retrieval frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); 2) improving the representation of external knowledge to facilitate retrieval and reasoning, with graph-based RAGs being a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
@@
Through a combination of theoretical analysis and empirical evaluation, we rethink the limitations of current RAG systems by introducing a novel perspective based on context recall and precision metrics. Our findings reveal that both regular and graph-based RAGs struggle not only to retrieve truly grounded context but also to accurately discern the relationship between retrieved content and the user query. We identify this fundamental issue as one primary reason why LLMs in RAG frameworks often generate seemingly relevant yet shallow responses that lack essential details.
To address these gaps, we introduce CausalRAG, a novel RAG framework that integrates causal graphs to enhance retrieval accuracy and reasoning performance. Unlike regular and graph-based RAGs, CausalRAG constructs a causal graph from uploaded documents, preserving contextual relationships while capturing cause-effect dependencies. By ensuring that retrieved documents are both relevant and causally grounded, CausalRAG enables the generation of more contextually rich and causally detailed responses. This approach not only improves retrieval effectiveness but also mitigates hallucinations and enhances answer faithfulness.
@@
We evaluate CausalRAG on datasets from diverse domains and across varying context lengths, comparing its performance with regular RAG and GraphRAG, a graph-based RAG framework from Microsoft Edge et al. (2024). Our experiments assess performance across three key metrics: answer faithfulness, context recall, and context precision. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We systematically identify the inherent limitations of RAG’s retrieval process through analytical and experimental study. More importantly, we uncover one major reason why LLMs in RAG tend to generate superficial, generalized answers that lack the grounded details expected by users.
  • We propose CausalRAG, a framework that enhances both retrieval and generation quality by incorporating causality into the RAG, effectively addressing these limitations.
  • Our work further mitigates hallucination issues and significantly improves the interpretability of AI systems. We summarize key findings and insights in both retrieval and generation process, contributing to research in RAG.
@@ 2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research primarily advances RAG along two key dimensions: 1) improving retrieval efficiency and integration mechanisms; 2) enhancing the representation of external knowledge to facilitate reasoning and interpretability.
Optimizing retrieval flow and interaction. The first stream focuses on improving the interaction flow within the RAG system to enhance output quality. Approaches have introduced pre-retrieval, retrieval, and post-retrieval refinements to mitigate redundancy and computational overhead Wang et al. (2024). Modular RAG architectures further advance this by enabling iterative retrieval-generation cycles, allowing dynamic interactions between retrieval and content creation. For example, CAiRE-COVID Su et al. (2020) demonstrated the effectiveness of iterative retrieval in multi-document summarization, while some work Feng et al. (2023) extended this approach to multi-hop question answering. Recent innovations include METRAG Gan et al. (2024), which integrates LLM supervision to generate utility-driven retrieval processes, and RAFT Zhang et al. (2024a), which trains models to disregard distractor documents while improving citation accuracy through chain-of-thought reasoning.
Structuring external knowledge for efficiency. The second direction explores improved methods for structuring external knowledge to achieve better retrieval efficiency. For example, GraphRAG Edge et al. (2024), proposed by Microsoft, treats external knowledge as interconnected nodes, capturing causal and thematic relationships to enhance retrieval depth and reasoning. LightRAG introduced a dual-level retrieval mechanism for incremental knowledge updates Guo et al. (2024). More recently, Lazy GraphRAG has been developed as a continuous version of GraphRAG, deferring computationally expensive operations until query time and leveraging lightweight indexing techniques for efficiency Potts (2024). Despite these advancements, ensuring the quality and relevance of retrieved documents remains a significant concern in current RAG systems, as it directly impacts the coherence of the final response Gupta et al. (2024).
@@ 2.2 Causal Graphs and RAG
The combination of causal graphs and RAG has emerged as a promising approach to enhancing knowledge retrieval and reasoning. As causality provides a structural understanding of dependencies within data, it enables more interpretable and reliable AI outputs Ma (2024). Existing research in this domain primarily advances causal discovery with RAG and LLMs. For instance, some work proposed an LLM-assisted breadth-first search (BFS) method for full causal graph discovery, significantly reducing time complexity Jiralerspong et al. (2024). Additionally, some further introduced a correlation-to-causation inference (Corr2Cause) task to evaluate LLMs’ ability to infer causation from correlation, revealing their limitations in generalization across different datasets (Jin et al., 2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One approach integrates causal graphs within the LLM architecture itself, structuring the transformer’s internal token processing using causality rather than enhancing RAG retrieval Chen et al. (2024b). Another employs causal graphs in RAG systems but focuses on the pre-retrieval stage and largely reduces the core process into a single embedding model without deeper exploration Samarajeewa et al. (2024). GraphRAG is a well-regarded and influential work in this area, as it introduces a graph-based structure into the RAG system, leveraging graph community detection and summarization techniques for retrieval Edge et al. (2024). While it does not incorporate causality, it has significantly improved RAG performance. Therefore, we adopt GraphRAG as the baseline for our work.
In the following sections, we first analyze the nature of regular RAG and grah-based RAGs from a novel perspective, identifying their inherent limitations. We then introduce a causal graph structure to address these gaps within the RAG system and present our framework—CausalRAG.
@@ 3.1 Limitations of regular RAG
Here $R$ is the reference set of all relevant reference documents. $C_{i}$ is the $i^{th}$ retrieved document. $\mathbb{I}(C_{i}\in R)$ is an indicator function that returns 1 if $C_{i}$ belongs to the reference set $R$, otherwise 0. It should always return 1 if no hallucination occurs in the LLM.
@@
Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to user’s query, otherwise 0.
Recall-Precision Perspective. Context recall measures how much of the correct contextual information can be retrieved from the user’s uploaded documents given a query. While RAG primarily focuses on retrieving related documents, adjusting retrieval parameters to increase the number of retrieved documents can generally lead to higher recall in reality. However, context precision presents a challenge. It measures the proportion of retrieved documents that are actually correct given the user query. As discussed earlier, RAG’s reliance on semantic similarity rather than causal relevance often leads to the retrieval of superficially similar but logically irrelevant content. In summary, while RAG can recall numerous answers from reference materials, the proportion of correct responses remains low, ultimately reducing its precision. This recall-precision perspective provides a new lens to see limitations of regular RAG.
@@ 3.2 Rethinking Graph-based RAGs
Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG. By summarizing and ranking the importance of graph communities before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision. However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information. More critically, its reliance on community-based summarization for retrieval may adversely impact recall. Based on these analytical insights, we further conducted an experimental study to validate our analysis.
Experimental study. As shown in Figure 1(b), we conducted an experimental study to empirically verify our analysis. Specifically, we employed an LLM to evaluate the performance of regular RAG and GraphRAG using the RAG evaluation framework Ragas Es et al. (2023). It takes as input the three key elements—user query, retrieved context, and reference—along with predefined metric definitions, and prompts the LLM to assign numerical ratings. This LLM-based evaluation has been widely used in recent RAG research and has also been adopted in RAG evaluations Samarajeewa et al. (2024); Edge et al. (2024).
The results indicate that both global and local versions of GraphRAG achieve higher context precision compared to regular RAG. However, its recall performance remains unsatisfactory, even slightly lower than that of RAG, due to its graph community-based retrieval process.
By combining our precision-recall-based analytical analysis with empirical findings, we clearly outline the inherent limitations of both RAG and its graph-based extensions. In the next section, we introduce our proposed framework, CausalRAG, designed to address these issues.
@@ 4 Methodology
In this section, we introduce our novel framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems. Overall, CausalRAG utilizes a graph-based approach to represent uploaded documents, enhancing robustness against discontinuities caused by text chunking (as shown in Figure 2). More importantly, by expanding and tracing nodes in the graph through causal relationships, CausalRAG is able to retrieve a more diverse set of contextual information while maintaining causal groundedness. This enables CausalRAG to achieve both high recall and strong precision. We now discuss each step in detail.
@@ 4.1 Indexing
At the outset, once the system receives the user’s uploaded documents and query, we first index these inputs into our vector database. For the uploaded documents, we employ a text-based graph construction method to transform the text into a structured graph before storing the nodes and edges in the vector database. Specifically, we leverage an LLM to construct the graph following the approach by LangChain Chase (2022), where the LLM scans the text to identify graph nodes and determine relationships between them. Although LLM-based graph construction has shown strong performance Chen et al. (2024a) and is widely adopted in graph-based RAG research , we further validate our constructed graph using expert knowledge, which we elaborate on in the case study. After constructing and indexing the base graph, we embed the user query into the vector database, preparing it for subsequent search and matching. Notably, this indexing process occurs independently of query time, allowing for efficient retrieval during inference.
@@ 4.2 Discovering and Estimating Causal Paths
This is the first step during query time. We begin by matching the user query to nodes in the graph based on their embedding distance. The $k$ nodes with the smallest distances are selected, representing the most relevant information directly related to the query. Notably, $k$ is a tunable parameter—higher values retrieve more relevant information at the cost of increased computational complexity.
After selecting the initial $k$ nodes, we expand the search along the base graph’s edges by a step size of $s$, thereby broadening the retrieved context. This step is crucial as it preserves causal and relational connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ controls the depth of expansion, where higher values lead to more diverse information retrieval.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph. As discussed, LLMs have demonstrated efficiency in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024). This step ensures that CausalRAG prioritizes causally relevant information, improving precision.
Furthermore, the derived causal graph serves two key purposes: 1) It preserves causally relevant information that traditional retrieval methods struggle to capture. More importantly, by adjusting the parameter $s$, this approach can capture long-range causal relationships within the text, particularly when the text is lengthy; 2) It filters out semantically related but causally irrelevant information. Without this filtering, responses may contain unnecessary or even hallucinated content, compromising answer faithfulness.
@@ 4.3 Retrieving Context Causally
After constructing the causal graph, we summarize the retrieved information and generate a causal summary. Notably, the input at this stage consists of information that is not only highly relevant but also causally grounded in the user’s query, ensuring greater validity. This approach contrasts with traditional retrieval methods, which often rely purely on semantic similarity and may retrieve contextually related yet causally irrelevant information.
The causal summary is derived by tracing key causal paths within the graph, prioritizing nodes and relationships that contribute directly to answering the query. This ensures that the retrieved information maintains logical coherence and factual consistency while filtering out spurious or weakly related context. Additionally, by leveraging causal dependencies, our method reduces the risk of retrieving semantically similar but misleading evidence.
Once the causal summary is generated, it is combined with the user query to construct a refined prompt for CausalRAG. This structured final input allows RAG to focus on reasoning through causal relationships rather than merely aggregating loosely related text spans.
@@ 5 Experiment
To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and GraphRAG baselines. Our evaluation spans multiple datasets, retrieval settings, and performance metrics, ensuring a comprehensive analysis of retrieval quality and answer faithfulness. We systematically explore different parameter settings and retrieval strategies to assess their impact on model performance.

### v2

@@ 1 Introduction
Large language models (LLMs) have transformed natural language processing (NLP), enabling a wide range of applications Anthropic (2024); Google (2024); OpenAI (2024). However, their reliance on static, pre-trained knowledge limits their ability to incorporate and reason over dynamically updated external information, particularly in knowledge-intensive domains. Retrieval-Augmented Generation (RAG) addresses this limitation by combining external retrieval with generative modeling to improve contextual understanding and response quality Lewis et al. (2021).
@@
Recent efforts to improve RAG focus on two fronts: 1) enhancing retrieval efficiency through adaptive and modular frameworks Gan et al. (2024); Ravuru et al. (2024); Zhang et al. (2024a); and 2) better structuring external knowledge, with graph-based RAGs emerging as a dominant approach Edge et al. (2024); Guo et al. (2024); Potts (2024). Despite these advancements, existing RAG architectures still face critical limitations that impact retrieval quality and response accuracy, primarily due to three key issues: 1) disruption of contextual integrity caused by the text chunking design; 2) reliance on semantic similarity rather than causal relevance for retrieval; and 3) a lack of accuracy in selecting truly relevant documents.
@@
Through theoretical and empirical analysis, we rethink the limitations of current RAG systems by introducing a novel perspective grounded in context recall and precision metrics. Using this lens, we find that both regular and graph-based RAGs often fail to retrieve causally grounded content or accurately align it with the user query. We identify this fundamental issue as a primary reason why LLMs in RAG frameworks frequently produce seemingly relevant yet shallow responses that lack essential details.
To address these gaps, we propose CausalRAG, a novel RAG framework that integrates causal graphs to guide retrieval. By identifying cause–effect relationships within external knowledge, CausalRAG preserves contextual coherence and improves reasoning fidelity. This leads to more accurate and causally grounded responses, while reducing hallucinations and enhancing answer faithfulness.
@@
We evaluate CausalRAG across datasets and varying context lengths, comparing it with regular and other competitive graph-based RAGs over multiple metrics. Results demonstrate that CausalRAG achieves superior performance across different contexts. Additionally, we conduct a case study and a parameter analysis to further examine our framework, analyzing and providing insights that contribute to ongoing research in RAG. The contributions of this work are threefold:
  • We identify the root limitations of current RAG systems through analytical and empirical studies, revealing why LLMs often produce superficial and under-grounded responses.
  • We propose CausalRAG, which improves retrieval and generation by grounding responses in causal structure.
  • Our work further mitigates hallucination issues and enhance interpretability, offering new insights into the design of RAG systems.
@@ 2.1 Retrieval-Augmented Generation
RAG enhances LLMs’ ability to handle knowledge-intensive tasks by integrating external knowledge retrieval Lewis et al. (2021). Existing research advances RAG in two main directions: 1) optimizing retrieval flow and interaction; and 2) structuring and utilizing external knowledge—particularly through knowledge graphs—for multi-stage reasoning.
Optimizing retrieval flow and interaction. The first stream focuses on refining retrieval dynamics to improve output quality. Several methods introduce pre-, mid-, and post-retrieval enhancements to reduce redundancy and computation Wang et al. (2024). Modular RAG architectures enable iterative retrieval–generation loops for more adaptive workflows. For instance, CAiRE-COVID Su et al. (2020) demonstrated multi-document summarization with iterative retrieval, and others extended this to multi-hop QA Feng et al. (2023). Recent work like METRAG Gan et al. (2024) leverages LLM supervision for utility-driven retrieval, while RAFT Zhang et al. (2024a) trains models to disregard distractor documents through chain-of-thought reasoning.
Structuring and utilizing external knowledge for multi-stage reasoning. The second stream focuses on representing and leveraging external knowledge—often via knowledge graphs—to support multi-level reasoning. GraphRAG Edge et al. (2024) represents documents as nodes and edges and summarizes subgraph communities for retrieval. LightRAG Guo et al. (2024) introduces dual-level retrieval and supports graph dynamic updates, while Lazy GraphRAG Potts (2024) further improves efficiency by delaying heavy computations until query time.
More recent multi-stage RAGs formalize hierarchical reasoning across multiple stages. PolyRAG Chen et al. (2025) uses a knowledge pyramid for coarse-to-fine retrieval. KG2RAG Zhu et al. (2025) guides RAG with KG-guided chunk expansion process to enhance factual grounding and control. GenTKGQA Gao et al. (2024) performs two-stage retrieval and generation over temporal subgraphs. HippoRAG2 Gutiérrez et al. (2025) integrates PageRank ranking with deeper passage reasoning for better answer quality. However, most of these approaches focus primarily on matching and ranking information within the knowledge graph, with limited attention to the causal intent behind the user’s query. A key challenge in RAG remains ensuring that retrieved information is not only relevant but also coherently aligned with the user’s underlying reasoning needs and the generation objective Gupta et al. (2024).
@@ 2.2 Causal Graphs and RAG
Causal graph integration has emerged as a promising way to improve retrieval and reasoning. Causal structures support more interpretable and reliable outputs by modeling underlying dependencies Ma (2024). Prior work in this area primarily focuses on causal discovery using RAG and LLMs. For instance, LLM-assisted BFS has been used to accelerate full causal graph discovery Jiralerspong et al. (2024), while Corr2Cause evaluates LLMs’ ability to infer causation from correlation Jin et al. (2024).
Despite these advancements, most studies focus on utilizing RAG or LLMs for causal discovery or causal effect estimation Ma (2024); Kıcıman et al. (2024), whereas the direct integration of causal graphs into RAG architectures remains largely unexplored. Our work aims to be a pioneer in this direction. A few existing studies have touched upon this concept but differ in scope. One integrates causality within the transformer architecture itself Chen et al. (2024), and another uses causal graphs only in pre-retrieval stages without deeper exploration Samarajeewa et al. (2024). GraphRAG Edge et al. (2024), while influential for its use of community detection and graph summarization, does not incorporate causality—nor do many recent multi-stage graph-based RAG frameworks.
In the following sections, we first analyze the limitations of regular and graph-based RAGs using a new recall–precision lens. We then introduce a causal graph structure to address these gaps.
@@ 3.1 Limitations of Regular RAG
Here $R$ is the reference set of all relevant references. $C_{i}$ is the $i^{th}$ retrieved reference. $\mathbb{I}(C_{i}\in R)$ is an indicator function that returns 1 if $C_{i}$ belongs to the reference set $R$, otherwise 0. It should always return 1 if no hallucination occurs in the LLM.
@@
Here $\mathbb{I_{Q}}(C_{i}\in R)$ is an indicator function that returns 1 if the context retrieved is causally related to the user’s query, otherwise 0.
Recall–precision perspective. Context recall measures the extent to which relevant contextual information can be retrieved from external knowledge given a query. In practice, increasing the number of retrieved documents typically improves recall in RAG systems. However, this semantic search–based approach often sacrifices precision—the proportion of retrieved content that is truly correct and directly relevant to the user query. For example, when asking, “How does this article define AI?”, regular RAG tends to retrieve all cited definitions that are semantically similar, even though only the author’s own definition is actually pertinent. This illustrates a core limitation of relying on semantic similarity rather than causal relevance: the retrieval of content that appears related but is logically irrelevant. More critically, low retrieval precision introduces systematic bias, diluting the accuracy and reliability of the retrieved context.
In summary, while RAG can recall numerous answers from reference materials, the proportion of correct context remains low, ultimately reducing its precision. This recall-precision perspective provides a new lens to see the limitations of RAG frameworks.
@@ 3.2 Rethinking Graph-based RAGs
Applying this perspective, we can better understand why Graph-based RAGs serve as improved variants of RAG. By summarizing and ranking the importance of graph information before retrieval, they largely enhance the quality of retrieved context, thereby improving context precision. However, it only partially addresses the identified limitations, as its summarization process does not entirely filter out irrelevant information. More importantly, its reliance on subgraph summarization for retrieval may adversely impact recall by omitting less graph-central but still causally relevant context. To further examine these trade-offs, we conducted an experimental study to empirically validate these analytical insights.
Experimental study. Figure 1(b) reports experimental results for Regular RAG and both variants of GraphRAG. The graph-based approaches markedly improve context precision, reflecting the additional reranking and subgraph-summary steps in their retrieval pipelines. This gain comes with a modest reduction in context recall, indicating that some relevant passages are pruned during summarization. These scores were obtained with the Ragas evaluation framework Es et al. (2023); more comprehensive experiments and implementation details are shown in Section 5.
By combining our recall–precision-based analytical insights with empirical findings, we highlight the inherent limitations of both standard RAG and its graph-based extensions. Specifically, we show that relying on semantic similarity and subgraph summarization—rather than causal relationships—introduces trade-offs that often result in superficial and less accurate generation. We also present a case study in Section 5.3, which explicitly compares the retrieval processes across different RAG frameworks. In the following section, we introduce our proposed framework, CausalRAG, developed to address these limitations through causally grounded retrieval.
@@ 4 Methodology
In this section, we introduce our proposed framework—CausalRAG—which integrates RAG with causality to overcome the limitations of existing RAG systems. Overall, CausalRAG constructs a text-based graph from uploaded documents and discovers causal paths among nodes to guide retrieval (Figure 2). By embedding the user query, matching relevant nodes, and expanding them along causally linked paths, the system generates a causal summary that serves as the retrieval context. This approach enables the framework to preserve contextual coherence and retrieve more targeted and causally relevant evidence. We now describe each step of the framework in detail.
@@ 4.1 Indexing
At the outset, upon receiving the user’s uploaded documents and query, the system first indexes both inputs into a vector database. For the documents, we employ a text-based graph construction method that transforms unstructured text into a structured graph comprising nodes and edges. Specifically, we follow the approach proposed in LangChain Chase (2022), where an LLM parses the text to identify key entities or concepts as nodes and infers relationships between them as edges. Though this is a widely adopted approach in RAG research, we further validate the resulting graphs using expert knowledge, as discussed in the case study (Section 5.3). Once the graph is constructed, it is embedded and stored in the vector database, enabling efficient similarity-based retrieval. The user query is also embedded at this stage, preparing the system for subsequent matching. Importantly, this indexing process is performed offline and independently of query time, ensuring fast and scalable inference.
@@ 4.2 Discovering and Estimating Causal Paths
At query time, we first match the user query to graph nodes using embedding distance and select the top-$k$ closest nodes. The parameter $k$ controls the retrieval breadth: higher values return more context but increase computational cost. We then expand these nodes along graph edges by a step size $s$, which allows the framework to preserve causally and relationally connections within the text, allowing CausalRAG to retrieve more context while maintaining high recall. The parameter $s$ determines the expansion depth and diversity of retrieved information.
Once the relevant nodes and edges are collected, we employ an LLM to identify and estimate causal paths within them, constructing a refined causal graph and generating the causal summary report (see LLM prompts in Appendix A). LLMs have demonstrated superiority in discerning and analyzing causal relationships Zhang et al. (2024b); Zhou et al. (2024) and this step ensures that CausalRAG prioritizes causally relevant information, improving precision. The resulting causal graph serves two purposes: (1) capturing long-range cause–effect dependencies often missed by semantic retrieval, particularly in longer texts; and (2) filtering out superficially related but causally irrelevant content, reducing hallucinations and improving faithfulness.
Life Sciences Computing & Math Social Sciences Physics Other Total Tokens
13.27% 14.29% 21.43% 14.29% 36.73% 21,285
Table 1: Statistics of the dataset domain distribution and token lengths.
Figure 3: Performance comparison of CausalRAG, regular RAG, and other graph-based RAGs across three key metrics: answer faithfulness, context recall, and context precision.
@@ 4.3 Retrieving Context Causally
After tracing and identifying key causal paths in the causal summary, we feed the result into the next stage and combine it with the user’s query (see LLM prompts in Appendix A). The causal summary prioritizes nodes that directly support the query, ensuring coherence and factual consistency while filtering out spurious information. This step reduces the risk of retrieving misleading yet superficially similar contexts. Finally, the combined prompt—including the causal summary—is used for generation, enabling CausalRAG to reason over major causal relationships rather than merely aggregating loosely related text spans.
@@ 5 Experiment
To evaluate the effectiveness of CausalRAG, we conduct a series of experiments comparing it with regular RAG and other competitive graph-based RAGs across multiple performance metrics to ensure a comprehensive assessment. In addition, we present a case study that explicitly compares the retrieval processes of different RAG variants, and a parameter study to further examine the behavior and performance of CausalRAG.